Corporate Fund started the year with a net asset value of $14.00. By year-end, its NAV equaled $13.20. The fund paid year-end distributions of income and capital gains of $2.20. 

What was the rate of return to an investor in the fund? (Round your answer to 2 decimal places.)

 Rate of return = _______ %
Answer #1

Rate of Return = (Final Value - Initial value + Distributions)/Initial value

Rate of Return = ($13.20 - $14 + $2.20)/$14

Rate of Return = 10.00%
Answer #2

Answer: 6.201% Explanation: Given that, Net asset value = $12.90 By year-end net asset value = $12.30 Fund paid year-end distributions of income and capital gains = $1.40 change in NAV: = By year-end net asset value - Net asset value =  $12.30 - $12.90 = -$0.6 Rate of Return: = (change in net asset value + Distributions) ÷ Start of Year net asset value = ( -$0.6 + $1.40) ÷ $12.90 = 0.8 ÷  $12.90 = 0.06201 or 6.201% 

Read more at Answer.Ya.Guru – https://answer.ya.guru/questions/5070978-corporate-fund-started-the-year-with-a-net-asset-value-of-1290.html
Answer #3

SOLUTION :


Gain in NAV in one year = 13.20 - 14 = - 0.80 ($)

 (negative value indicates that there is loss)


Income distribution and CG = 2.20 ($)


Net gains = 2.20 - 0.80 = 1.40 ($)


Rate of return to an investor in the fund 

= Net Gains / Initial NAV *100 (%)

= 1.40 / 14 *100. (%)

= 10% (ANSWER)


answered by: Tulsiram Garg